The cost of new machine will be $100000 after 12 years. In that time the old machine can be sold for $5000 as salvage value. What amount should be retained out of profits at the end of each year accumulate at C.I & P.A?
1
Expert's answer
2020-09-21T15:55:49-0400

To get the amount to be retained out of profits per year ;


="\frac{Cost of Asset - Salvage Value} 
{Useful Life of Asset}"


="(\frac{100,000-5,000}{12})"


="\frac{95,000}{12}"


=$ 7,916.67